### v2

# Ergotropic advantage in a measurement-fueled quantum heat engine

Abstract
This paper investigates a coupled two-qubits heat engine fueled by generalized measurements of the spin components and using a single heat reservoir as sink. Our model extends the proposal of Yi and coworkers [Phys. Rev. E 96, 022108 (2017)] where the role of a hot reservoir in a four-stroke cycle was replaced by a quantum measurement apparatus, the other steps being two quantum adiabatic strokes and thermalization with a cold reservoir. We propose a five-stroke cycle, where an ergotropy extracting stroke is introduced following the measurement stroke, and study the effect of measurements of different spin components on the performance of the machine. For measurements along z-z directions, we find two possible occupation distributions that yield an active state and the ergotropic stroke improves the performance of the engine over the four-stroke cycle. Further, the three-stroke engine ( without the adiabatic strokes) yields the same performance as the five-stroke engine. For arbitrary working medium and non-selective measurements, we prove that the total work output of a five-stroke engine is equal to the sum of the work outputs of the corresponding four-stroke and three-stroke engines. For measurement directions other than z-z, there may be many possible orderings of the post-measurement probabilities that yield an active state. However, as we illustrate, for specific cases (e.g. x-x), a definite ordering may be obtained with the projective measurements. Thus, we find that the five-stroke engine exploiting ergotropy outperforms both its four-stroke as well as three-stroke counterparts.
I Introduction
Quantum thermodynamics [1, 2, 3, 4, 5, 6, 7] involves, amongst other pursuits, the study of thermal machines based on a quantum working medium. The idea of a quantum heat engine was first proposed by Scovil and Schulz-DuBois [8], who argued that a three-level maser, in simultaneous contact with a hot and a cold reservoir, can be regarded as a heat engine. State-of-the-art technology now allows us to experimentally control systems at the quantum scale, such as spins [9, 10, 11, 12, 13, 14, 15], superconducting qubits [16, 17] and trapped ions [18, 19]. So, the non-classical resources such as entanglement, quantum coherence, quantum correlations and noise are being hotly pursued for their use in improving the performance of thermal machines away from the classical regime [20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
In recent years, quantum measurements [30, 31] have attracted increasing attention due to their ability to mediate energy exchange—either as work extraction [32] or as effective heat transfer. Measurement-induced energy injection can mimic the role of a hot reservoir, provided the measured observable does not commute with the system Hamiltonian [33, 34, 35, 36]. Importantly, such engines operate without feedback control, distinguishing them from Szilard-type engines [37, 38, 39]. Related developments include quantum refrigerators and cooling protocols [40, 41], weak and continuous measurement schemes [42, 43, 44], non-ideal measurements [45] and work-fluctuation analyses [46]. Measurements further induce entropy production, probe–system correlations, and raise questions about associated energetic costs [47, 48, 49]. The classification of measurement-induced energy exchange and its role as a fueling mechanism [50, 51] remains an active research direction.
The main idea of our investigation is that after the non-selective quantum measurement injects heat into the working medium, the latter may be left in a non-passive or active state. From this state, it is possible to extract work by a unitary process in which the Hamiltonian of the system undergoes a cyclic evolution and the maximum work so extracted is known as the ergotropy [52]. The present work extends the measurement-fueled engine [33, 34, 53] by using weak quantum measurements on a coupled-qubits system while exploiting the active nature of the post-measurement state of the system. We incorporate an ergotropy-extracting stroke [54, 55, 56, 57, 58] after the measurement stroke, thereby upgrading the four-stroke cycle to a five-stroke cycle. With measurements on the spin components along specific directions, we highlight the thermodynamic advantage of the ergotropic stroke by comparing the performance of the five-stroke engine with the four-stroke (without ergotropy) as well as a three-stroke (without adiabatic strokes) cycle.
The paper is organized as follows. Section II describes the five-stroke heat cycle of the measurement-fueled engine, which is applied to a coupled qubits working medium with z-z measurements in Section III.1, while comparing with the four-stroke and the three-stroke cycles in Section III.2 and III.3 respectively. Section IV describes the results for other measurement directions. We conclude our analysis in Section V.
II Five-stroke engine
The working medium, or the system, is described by the Hamiltonian $H(B)$, with its initial state as the thermal state ($\rho^{\rm in}$) corresponding to a heat reservoir at inverse temperature $\beta$, and the control parameter set at $B=B_{2}$. We follow the convention of the two-reservoir model in which the (magnetic field) parameter at the cold reservoir is labelled as $B_{2}$ and in the present case, the only available reservoir acts as a cold reservoir. The occupation probability for the $n^{\rm th}$ energy eigenstate is labelled as $p_{n}$. The system undergoes a heat cycle consisting of five strokes, as described below (see Fig. 1).
Figure 1: A five-stroke engine cycle consisting of the first quantum adiabatic stroke ($1\rightarrow 2$), the measurement stroke ($2\rightarrow 3$), the ergotropy extraction stroke ($3\rightarrow 4$), the second quantum adiabatic stroke ($4\rightarrow 5$) and finally the thermalization stroke with the heat reservoir ($5\rightarrow 1$). Without the stroke $3\rightarrow 4$, the cycle reduces to the four-stroke engine of Ref. [33].
Stroke $1\to 2$: This is a quantum adiabatic process (AP-1) in which the parameter $B$ is increased from $B_{2}$ to $B_{1}$ sufficiently slowly such that the quantum adiabatic theorem [59] holds and the occupation probabilities remain unchanged. Since the system remains in its instantaneous energy eigenstate, the state at the end of this stroke is given by: $\rho=\sum_{n}p_{n}|n,B_{1}\rangle\langle n,B_{1}|$, where $|n,B_{1}\rangle$ are energy eigenstates at the end of this process.
Stroke $2\to 3$: Keeping the Hamiltonian fixed at $H(B_{1})$, a generalized discrete measurement of an observable $\cal O$} (incompatible with $H(B_{1})$) is made. This measurement changes the state of the system, with the post-measurement occupation probabilities denoted as $p^{\rm{\rm PM}}_{n}$.
Stroke $3\to 4$: This is termed as the ergotropic stroke. Following Ref. [52], the ergotropy is defined as the maximum work extractable from a quantum system via a cyclic unitary on its Hamiltonian. The state of the system from which work can be so extracted is denoted as an active state, otherwise it is called a passive state. Thus, the stroke $3\to 4$ is executed only if the post-measurement state is an active state (the criterion is explained in the next Section). In other words, if the post-measurement state is a passive state, then this stroke is not executed and the proposed heat cycle reduces to the four-stroke cycle of Ref. [33], with all the other steps remaining the same.
Stroke $4\to 5$: This is again a quantum adiabatic process (AP-2) in which the system is isolated and the parameter $B$ is slowly brought back to its initial value $B_{2}$.
Stroke $5\to 1$: Finally, the system thermalizes with the heat reservoir and returns to the initial state, thus completing the cycle.
The exchange of heat or work in a stroke is defined as the difference of the final and the initial mean energy of the system during that stroke. The work in the AP-1 stroke is given by
$W_{1}=\sum_{n}[E_{n}(B_{1})-E_{n}(B_{2})]p_{n}.$ (1)
[…]
$Q_{\rm M}=\sum_{n}E_{n}(B_{1})[p^{\rm{\rm PM}}_{n}-p_{n}]>0.$ (2)
To implement the ergotropic stroke, suppose that under the action of a cyclic unitary $U$, the post-measurement state $\rho^{\rm{PM}}$ transforms to $U\rho^{\rm PM}U^{{\dagger}}$ while the Hamiltonian returns to its original form $H(B_{1})$. The average work extracted in this process is given by $W(\rho^{\rm{PM}},U)={\rm Tr}[H(U\rho^{\rm{PM}}U^{{\dagger}}-\rho^{\rm{PM}})]<0$. The maximum of $W(\rho^{\rm{PM}},U)$ over all cyclic unitaries is called the ergotropy $W_{\rm erg}$. After the extraction of ergotropy (stroke $3\to 4$), the passive state is given by $|\rho^{\prime}=\sum_{n}p_{n}^{\prime}\ket{n,B_{1}}\bra{n,B_{1}}$, where due to the unitary nature of the ergotropic stroke, $p_{n}^{\prime}$ are just the shuffled eigenvalues of $\rho^{\rm PM}$ such that $p_{n}^{\prime}>p_{n+1}^{\prime}$ for $E_{n}(B_{1})<E_{n+1}(B_{1})$. Then, the ergotropy is given by
$W_{\rm{\rm erg}}=\sum_{n}E_{n}(B_{1})[p^{\prime}_{n}-p^{\rm{PM}}_{n}]<0.$ (3)
Next, the work in the AP-2 stroke is
$W_{2}=\sum_{n}[E_{n}(B_{2})-E_{n}(B_{1})]p_{n}^{\prime}.$ (4)
Finally, the heat rejected to the reservoir is
$Q_{\rm res}=\sum_{n}E_{n}(B_{2})[p_{n}-p^{\prime}_{n}]<0.$ (5)
The sign of $Q_{\rm res}$ may be argued on thermodynamic grounds as follows. The state of the working medium or the system undergoes a cycle. The measurement step causes an increase in the entropy of the system while the adiabatic strokes keep the entropy unchanged. So, in order to return to initial state, the system must release heat/entropy to the reservoir, implying that $Q_{\rm res}<0$.
Using energy conservation, the total work extracted in a cycle is
$W_{\rm T}^{(5)}=-W_{1}-W_{2}-W_{\rm erg}=Q_{\rm M}+Q_{\rm res}.$ (6)
The operation of the engine requires $W_{\rm T}^{(5)}>0$. So, the efficiency is defined as
$\eta=\frac{W_{\rm T}^{(5)}}{Q_{\rm M}}.$ (7)
In the next section, we apply the above framework to a coupled qubits working medium where we consider spin measurements on each qubit, along specific directions.
III Coupled-qubits working medium
We consider two coupled qubits following a 1D Hamiltonian with isotropic Heisenberg interaction [60]:
$H(B)=B(\sigma_{z}^{\rm A}\otimes I^{\rm B}+I^{\rm A}\otimes\sigma_{z}^{\rm B})+2J\!\sum_{i=x,y,z}\sigma_{i}^{\rm A}\otimes\sigma_{i}^{\rm B}.$ (8)
[…]
Now, the initial state is $\rho^{\rm{in}}={e^{-\beta H(B_{2})}}/{Z}$, where $Z=\sum_{n=1}^{4}e^{-\beta E_{n}}$ is the partition function. Then, we have $p_{n}=e^{-\beta E_{n}}/Z$. In the first adiabatic stroke, the work [Eq. (1)] is evaluated to be
$W_{1}=$ $-2(B_{1}-B_{2})\left(p_{2}-p_{4}\right)<0.$ (9)
We choose $M_{\alpha}$ operators in arbitrary directions $\hat{n}^{\rm A}$ and $\hat{n}^{\rm B}$, in the following form:
$\begin{split}M_{\pm,\pm}&=(c_{0}I^{\rm A}\pm c_{1}\vec{\sigma}^{\rm A}.\hat{n}^{\rm A})\otimes(c_{0}I^{\rm B}\pm c_{1}\vec{\sigma}^{\rm B}.\hat{n}^{\rm B})\\ M_{\pm,\mp}&=(c_{0}I^{\rm A}\pm c_{1}\vec{\sigma}^{\rm A}.\hat{n}^{\rm A})\otimes(c_{0}I^{\rm B}\mp c_{1}\vec{\sigma}^{\rm B}.\hat{n}^{\rm B})\end{split}$ (10)
where $c_{0}$ and $c_{1}$ are real parameters. From the completeness relation: $2c_{0}^{2}+2c_{1}^{2}=1$. One of the parameters, say $c_{0}$, can be chosen to define the strength of the measurement. Thus, $c_{0}=c_{1}=1/2$ implies a strong or projective measurement [34]. The general expressions for the post-measurement probabilities are quite complicated and it is not feasible to determine the conditions for their relative ordering. Now, we study measurements along specific directions.
III.1 z-z measurement
We choose $\vec{\sigma}^{\rm A}.\hat{n}^{\rm A}=\sigma_{z}^{\rm A}$ and $\vec{\sigma}^{\rm B}.\hat{n}^{\rm B}=\sigma_{z}^{\rm B}$. The post-measurement occupation probabilities are given by
$\begin{split}p_{1}^{\rm{PM}}&=p_{1}-4c_{0}^{2}(1-2c_{0}^{2})(p_{1}-p_{3}),\\ p_{2}^{\rm{PM}}&=p_{2},\\ p_{3}^{\rm{PM}}&=p_{3}+4c_{0}^{2}(1-2c_{0}^{2})(p_{1}-p_{3}),\\ p_{4}^{\rm{PM}}&=p_{4}.\end{split}$ (11)
The heat absorbed by the system during the measurement stroke [Eq. (2)] is
$Q_{\rm M}=8J(p^{\rm{PM}}_{3}-p_{3})>0.$ (12)
As a result of the measurement, the occupation probabilities of levels $E_{2}$ and $E_{4}$ are unchanged, while those of $E_{1}$ and $E_{3}$ come closer to each other. Intuitively, we see that the post-measurement probability distribution becomes more uniform and so its Shannon entropy increases. Also, $p^{\rm{PM}}_{1}-p^{\rm{PM}}_{3}=(4c_{0}^{2}-1)^{2}(p_{1}-p_{3})\geq 0$, where the equality is obtained for $c_{0}=1/2$. Note that $p_{4}$ is still the lowest probability in the distribution (11) since the probabilities were initially ordered as $p_{1}>p_{2}>p_{3}>p_{4}$. Based on these considerations, we can infer: $p^{\rm{PM}}_{1}\geq p^{\rm{PM}}_{3}>p_{4}$. Note that $\rho^{\rm PM}$ has no coherence in the energy eigenbasis for the z-z case. Therefore, to ascertain the active nature of the state $\rho^{\rm{PM}}$, we have to determine the relative ordering of the probabilities $p^{\rm{PM}}_{n}$. It can be seen that an active post-measurement state implies one of the following two possibilities:
$({\rm R1})\qquad$ $p^{\rm{PM}}_{1}\geq p^{\rm{PM}}_{3}>p_{2}>p_{4},$
$({\rm R2})\qquad$ $p_{2}>p^{\rm{PM}}_{1}\geq p^{\rm{PM}}_{3}>p_{4}.$
Figure 2: R1 case: Heat and work contributions in the five-stroke cycle vs. $c_{0}$. Parameters are set at $B_{1}=3.5$, $B_{2}$ = 3, $J$ = 1 and $\beta$ = 1.
Let us consider the R1 case in detail. Now, in the ergotropy stroke, the occupation probabilities $p^{\rm PM}_{n}$ get reordered in an opposite sense to the energy eigenvalues [52], implying that $p_{1}^{\prime}=p^{\rm{PM}}_{1}$, $p_{2}^{\prime}=p^{\rm{PM}}_{3}$, $p_{3}^{\prime}=p_{2}$ and $p_{4}^{\prime}=p_{4}$. Thus, the system state after the ergotropy stroke can be given as
[…]
$|.\displaystyle+p_{4}\ket{00}\bra{00}$. (13)
Note that R1 case is characterized by the condition $p^{\rm{PM}}_{3}>p_{2}$, which can be cast in the form:
$4c_{0}^{2}(1-2c^{2}_{0})>\frac{e^{2\beta B_{2}}-1}{e^{8\beta J}-1}=\chi(B_{2},J,\beta),$ (14)
which implies that R1 case bounds the parameter $c_{0}$ as follows: $(1-\sqrt{1-2\chi})/4<c_{0}^{2}<(1+\sqrt{1-2\chi})/4$, where $\chi<1/2$.
Next, the ergotropy [Eq. (3)] is given by
$W_{\rm{erg}}=-2B_{1}(p^{\rm{PM}}_{3}-p_{2})<0.$ (15)
Work in the second adiabatic stroke [Eq. (4)] is
$W_{2}$ $=2(B_{1}-B_{2})\left(p^{\rm{PM}}_{3}-p_{4}\right)>0.$ (16)
Finally, the heat exchange in the thermalization stroke [Eq. (5)] is
$Q_{\rm res}=2B_{2}(p^{\rm{PM}}_{3}-p_{2})-8J(p^{\rm{PM}}_{3}-p_{3}).$ (17)
Using Eq. (11), we can show that $Q_{\rm res}<0$, as required. The net work extracted is
$W_{\rm T}^{(5)}=Q_{\rm M}+Q_{\rm res}=2B_{2}\left(p^{\rm{PM}}_{3}-p_{2}\right)>0.$ (18)
In Fig. 2, the above heat and work contributions are plotted for a specific example. It is observed that the work output becomes optimal at $c_{0}$ = 1/2 i.e. when $p^{\rm{PM}}_{3}$ achieves its maximum value $(p_{1}+p_{3})/2$. The efficiency in the R1 case can be expressed as
$\eta=$ $\frac{B_{2}}{4J}\left[1-\frac{(e^{2\beta B_{2}}-1)}{4c_{0}^{2}(1-2c_{0}^{2})\left(e^{8\beta J}-1\right)}\right].$ (19)
It is interesting to note that $\eta$ is independent of the parameter $B_{1}$. This is not a generic feature, but depends on the energy spectrum. For the given model, $\eta$ is maximum at $c_{0}=1/2$ i.e. for projective measurements. The efficiency is plotted in Fig. 3. For a given $c_{0}$, $\eta$ increases with $\beta$. For low enough temperatures, $\eta$ grows rapidly with $c_{0}$ and approaches its limiting value $B_{2}/4J$. In Fig. 4, the efficiency is plotted versus $J$ for a fixed $c_{0}$ at various temperatures. Again, decreasing the temperature increases the efficiency at low $J$ values, while for high $J$ values, all efficiency curves merge into each other and goes to zero.
Similarly, we can study the R2 case which is characterized by the ordering: $p_{2}>p^{\rm{PM}}_{1}\geq p^{\rm{PM}}_{3}>p_{4}$. To appreciate how this ordering may come about, consider the difference: $p_{2}-p^{\rm{PM}}_{1}=\left(p_{2}-p_{1}\right)+4c_{0}^{2}(1-2c^{2}_{0})\left(p_{1}-p_{3}\right)$. As we are working in the regime $4J>B_{2}$, so $B_{2}$ close to $4J$ implies that the lowest two levels are close to each other. Then, the term $(p_{2}-p_{1})$ will be negligible and the condition $p_{2}>p^{\rm PM}_{1}$ can prevail, thereby producing the R2 case.
The total work output is
$W_{\rm T}^{(5)}=$ $2(4J-B_{2})\left(p_{2}-p^{\rm{PM}}_{1}\right)>0,$ (20)
whereas $Q_{\rm M}$ is given by Eq. (12). The efficiency of this case can then be computed. As shown in Fig. 5, the R2 case yields a rather small efficiency. Another notable feature is that upon decreasing the temperature, the efficiency decreases, thus showing an opposite trend to the R1 case.
Figure 3: R1 case: Efficiency [Eq. (19)] vs. $c_{0}\in[0,1/\sqrt{2}]$. $B_{2}=3$ and $J=1$. So, the limiting value of the efficiency (as $\beta\to\infty$) is $B_{2}/4J=0.75$.
Figure 4: R1 case: Efficiency [Eq. (19)] vs. $J$. $B_{2}=3$ and $c_{0}=0.3$.
III.2 Four-stroke cycle
For comparison, we also consider the four-stroke cycle [33, 34] that does not involve the ergotropic stroke. It implies the occupation probabilities in AP-2 are given by $p^{\rm{PM}}_{n}$. The expressions for $W_{2}$ and $Q_{\rm res}$ are obtained by replacing $p_{n}^{\prime}$ with $p^{{\rm PM}}_{n}$ in the corresponding expressions for the five-stroke cycle as in Sec. II. Thus, the total work extracted can be expressed as:
$W_{\rm T}^{(4)}=\sum_{n}[E_{n}(B_{2})-E_{n}(B_{1})](p_{n}-p_{n}^{\rm PM}).$ (21)
For the two-qubits system and with z-z measurements, we obtain that the net work output of the four-stroke cycle is zero. This feature was noted in Ref. [34] for z-z projective measurements and we affirm it for generalized measurements too. Thus, we can conclude that the ergotropy stroke improves the performance of the engine based on generalized z-z measurements.
Figure 5: R2 case: Efficiency vs. $c_{0}$. $B_{2}=3.9$, $J=1$. The efficiency decreases as the temperature is lowered at a given measurement strength, unlike the R1 case (see Fig. 3). The R2 efficiency is also maximized at $c_{0}=1/2$.
III.3 Three-stroke cycle
As an alternative, we consider an ergotropy based three-stroke cycle consisting of the following steps. We prepare the system with Hamiltonian $H(B_{2})$ in a thermal state $\rho^{\rm in}$ at inverse temperature $\beta$. i) The system is isolated from the reservoir (assuming a weak interaction with the reservoir and so negligible costs in attaching and detaching the system from the reservoir) and nonselective generalized measurements are performed which transform the state to $\rho^{\rm PM}$. Heat transferred due to measurement is given by
$Q_{\rm M}=\sum_{n}E_{n}(B_{2})[p^{\rm{\rm PM}}_{n}-p_{n}]>0.$ (22)
ii) Ergotropy, $W_{\rm erg}$, is extracted which yields the passive state $\rho^{\prime}$. iii) Finally, the system is brought to the initial state by thermalization with the reservoir. The net work extracted, equal in magnitude to the ergotropy in step (ii), is given by
$W_{\rm T}^{(3)}=-W_{\rm erg}=\sum_{n}E_{n}(B_{2})(p_{n}^{\rm PM}-p_{n}^{\prime}).$ (23)
From Eqs. (6), (21) and (23), we obtain an interesting equality:
$W_{\rm T}^{(5)}=W_{\rm T}^{(4)}+W_{\rm T}^{(3)}.$ (24)
Note that the above relation holds for arbitrary Hamiltonian and nonselective measurements. It implies that the performance of a four-stroke measurement engine ($W_{\rm T}^{(4)}\geq 0$), with an active post-measurement state, can be enhanced by using the five-stroke cycle.
For the R1 case, we obtain $W_{\rm T}^{(3)}=2B_{2}(p^{\rm{PM}}_{3}-p_{2})$, which, as expected, is equal to the work in the five-stroke cycle [Eq. (18)] since $W_{\rm T}^{(4)}=0$ for z-z measurements. $Q_{\rm M}=8J(p^{\rm{PM}}_{3}-p_{3})>0$ for the three-stroke cycle, which is the same as Eq. (12). Therefore, its efficiency is the same as for the five-stroke cycle [Eq. (19)] in the case of z-z measurements. A similar conclusion holds for the case R2.
IV Other measurement directions
We also studied the performance of the engine with generalized measurements along x-x (y-y), x-y, and x-z (y-z) directions. In general, the post-measurement state $\rho^{\rm PM}$ contains some coherence (in the energy basis), making it an active state. Then, to determine the passive state after the ergotropic stroke, we need to order the eigenvalues of the state $\rho^{\rm PM}$ in a non-increasing sense. To illustrate, we consider x-x measurements for which we set $\vec{\sigma}^{\rm A}.\hat{n}^{\rm A}=\sigma_{x}^{\rm A}$ and $\vec{\sigma}^{\rm B}.\hat{n}^{\rm B}=\sigma_{x}^{\rm B}$ in Eq. (10). For the case of x-x weak measurements, the state $\rho^{\rm PM}$ is active, and we get the post-measurement occupation probabilities as
[…]
The relative ordering of the above probabilities is not straightforward, though we can prove specific inequalities, such as $p^{\rm{PM}}_{1}>p^{\rm{PM}}_{3}$. For $c_{0}<1/2$, $p^{\rm{PM}}_{2}<p^{\rm{PM}}_{4}$ while $p^{\rm{PM}}_{2}>p^{\rm{PM}}_{4}$ for $c_{0}>1/2$. Unlike the case of z-z measurements, here the four-stroke cycle with weak measurements yields a non-zero total work.
The eigenvalues of $\rho^{\rm PM}$ can be written as
$\begin{split}p_{1}^{\prime}&=p^{\rm{PM}}_{1},\\ p_{2}^{\prime}&=\frac{p^{\rm{PM}}_{2}+p^{\rm{PM}}_{4}}{2}+\Delta,\\ p_{3}^{\prime}&=p^{\rm{PM}}_{3},\\ p_{4}^{\prime}&=\frac{p^{\rm{PM}}_{2}+p^{\rm{PM}}_{4}}{2}-\Delta,\end{split}$ (26)
where $\Delta>0$ depends in a complicated way on $p_{n}$ and $c_{0}$. The above eigenvalues are also the occupation probabilities in the passive state $\rho^{\prime}$, to be ordered in a decreasing sense. However, there can be different orderings of these probabilities in different parameter regimes. On the other hand, for the specific case of projective measurements ($c_{0}=1/2$), the post-measurement probabilities are simplified as: $p^{\rm{PM}}_{1}=(1+p_{1}-p_{3})/4,\;p^{\rm{PM}}_{2}=p^{\rm{PM}}_{4}=1/4,\;p^{\rm{PM}}_{3}=(1-p_{1}+p_{3})/4$, while the eigenvalues of $\rho^{\rm PM}$ are $p_{1}^{\prime}=p_{2}^{\prime}=p^{\rm{PM}}_{1}$ and $p_{3}^{\prime}=p_{4}^{\prime}=p^{\rm{PM}}_{3}$, which are already ordered i.e. $p_{1}^{\prime}=p_{2}^{\prime}>p_{4}^{\prime}=p_{3}^{\prime}$, since $p^{\rm{PM}}_{1}>p^{\rm{PM}}_{3}$. In this special case, the heat exchanged during the measurement stroke is calculated to be
$Q_{\rm M}=2B_{1}(p_{2}-p_{4})+2J(2p_{1}-p_{2}-p_{4}).$ (27)
The ergotropy is given by $W_{\rm erg}=-B_{1}(p_{1}-p_{3})<0$, while the total work output is
$W_{\rm T}^{(5)}=2(B_{1}-B_{2})(p_{2}-p_{4})+B_{2}(p_{1}-p_{3})>0.$ (28)
The work output of a four-stroke cycle with projective measurements is $W_{\rm T}^{(4)}=2(B_{1}-B_{2})(p_{2}-p_{4})<W_{\rm T}^{(5)}$, while the three-stroke cycle yields $W_{\rm T}^{(3)}={B_{2}}(p_{1}-p_{3})$. Thus, we verify the equality (24).
In Fig. 6, we show the effect of weak measurements as they can lead to a higher efficiency compared to projective measurements. Here, the ordering of the probabilities in the post-ergotropy state is performed numerically for the given choice of system parameters. Similarly, we find that the post-measurement state for the x-z case has coherence and so is an active state. On the other hand, with x-y projective measurements, we find that the post-measurement probability distribution is uniform ($p_{n}^{\rm PM}=1/4$). Being a passive state, the ergotropic stroke does not help in this case.
Before closing this section, we consider the case of spin-measurements along arbitrary directions. As mentioned earlier, ordering relations between the post-measurement probabilities or between the eigenvalues of $\rho^{\rm PM}$ are hard to obtain here. However, if we restrict ourselves to projective measurements, the following inequality always holds.
$p^{\rm{PM}}_{2}-p^{\rm{PM}}_{4}=$ $\frac{1}{2}(\cos^{2}{\theta_{\rm A}}+\cos^{2}{\theta_{\rm B}})(p_{2}-p_{4})\geq 0.$ (29)
The equality $p^{\rm{PM}}_{2}=p^{\rm{PM}}_{4}$ is recovered for projective measurements in $xy$-plane for each qubit ($\theta_{\rm A}=\theta_{\rm B}=\pi/2$) irrespective of the angles $\phi_{\rm A}$ and $\phi_{\rm B}$. On the other hand, for z-z measurements ($\theta_{\rm A}$ = $\theta_{\rm B}$ = 0), we obtain $p^{\rm{PM}}_{2}=p_{2}$ and $p^{\rm{PM}}_{4}=p_{4}$, as found in Section III.1.
Figure 6: x-x case: Efficiency vs. $J$ for the five-stroke engine. Weak measurements with $c_{0}<1/2$ can lead to a higher efficiency than with projective measurements ($c_{0}=1/2$). For comparison, the efficiency of the four-stroke engine with projective measurements, $\eta^{(4)}$, is also plotted. Here, $B_{1}=3.5$, $B_{2}=3$ and $\beta=1$.
V Conclusions
In this paper, we studied the effect of additional ergotropy extraction on the performance of a measurement-fueled quantum engine based on a coupled two-qubits system. Here, the weak measurements of the spin components are used as a resource for input heat. We first analyzed the case of z-z weak measurements in a five-stroke cycle. The corresponding four-stroke cycle yields zero work output. Notably, the post-measurement state has no coherence in the energy eigenbasis. However, there are two possible orderings of the post-measurement probabilities leading to an active post-measurement state. We find that an ergotropy-extracting stroke inserted after the measurement stroke enables a net positive work output. Furthermore, a reduced three-stroke cycle (without the adiabatic strokes) yields the same performance as the five-stroke cycle. However, for other measurement directions (e.g., x–x, x–z), the post-measurement active state, in general, has coherence. Here, the four-stroke engine already delivers non-zero work, yet the inclusion of an ergotropic stroke enhances the performance. We also find that weak x–x measurements can surpass the projective ones in efficiency, highlighting the role of the measurement strength as a tuning parameter. In contrast, some settings (e.g., x–y case with projective measurements) yield passive post-measurement states, offering no ergotropic advantage.
A general identity, valid for arbitrary Hamiltonian and non-selective measurements, was established, which shows that the total work of the five-stroke engine equals the sum of its four-stroke and three-stroke counterparts. The present findings underline three key points: (i) measurements combined with ergotropy extraction can increase useful work; (ii) engine performance depends sensitively on the measurement strength; and (iii) appropriate choices of these parameters can outperform both standard four-stroke and three-stroke cycles.
In our analysis, we have neglected some of the energetic and thermodynamic costs towards the implementation of the unitary protocols [62, 63]. The notion of work extraction from a quantum system assumes some coupling with a work reservoir, or a quantum battery. In certain situations, with a battery of a finite size and a finite coherence, the work extracted is expected be less than ergotropy [64]. One possibility is to assume that the battery is large enough and highly coherent, so that all the energy transferred to the battery can be interpreted as work. Further, the post-measurement state of the working medium may have some degree of coherence. Then, extracting the coherent part of the ergotropy also becomes an experimental challenge. We have assumed a quasi-static heat cycle. Thus, in the present framework, we take the costs incurred through these time-dependent protocols [65] to be negligible. However, in the corresponding finite-time thermodynamic cycle, these costs can be important, and accordingly, the comparison between the performance of four-stroke and five-stroke engines has to be reassessed.
In conclusion, we have demonstrated that combining measurement-induced population reshaping with the extraction of ergotropy provides a viable thermodynamic advantage for quantum engines. These features can be exploited in feasible experimental platforms such as NMR [51] and superconducting circuits [65] which have so far formulated state of the art protocols for single qubit cases. Future work can include many-body systems, feedback-based protocols and bounding the performance due to finite-time effects.
Acknowledgements.
S.J. thanks Sachin Sonkar for useful discussions and acknowledges financial support in the form of Senior Research Fellowship from the Indian Institute of Science education and Research Mohali.